# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

### v12

@@ 2.2 Tsukishima: A Representative Manufacturing Area
Finally, I provide a brief overview of the lives of the factory workers’ households in Tsukishima. Factories and workers’ residences were scattered throughout the island. The factory workers lived in similar tenements and used the commercial stores and restaurants in the central lot. [^17] Tsukishima is composed of Ishikawajima island at the northern end, the adjacent Tsukudajima and Shin-tsukudajima islands, and the rectangular Lots 1–3 in the south (Online Appendix Figure A.3). Specifically, factories and worker residences were scattered throughout Shin-tsukudajima and Lots 1 and 2, whereas Lot 3, a relatively new reclaimed site completed in 1913, was undeveloped. Online Appendix A.4 summarizes the housing in Tsukishima in detail using a housing survey. There were $209$, $109$, $74$, and $120$ stores selling daily necessities, clothing, tools, and food, drink, and entertainment, respectively. Regarding medical personnel, there were eight doctors, four dentists, eight obstetricians, and $14$ drug stores. There were also ten bathhouses (Department of Health, Ministry of the Interior 1923a, pp. 41–47). There were two postal offices, $13$ pawnshops, and five moneylenders, which were scattered throughout the island (Online Appendix Figure A.2). The main routes to the mainland were the wooden Aioi Bridge (Aioi bashi) and three ferries (watashi bune). The average daily total number of people using the ferries in 1918 was reported to be $53,318$, accounting for more than twice as many people as the Tsukishima population. [^18] See Department of Health, Ministry of the Interior (1923a, p. 30). The number of people and households in Tsukishima at the end of December 1918 was reported to be $24,399$ and $5,562$, respectively (National Police Agency 1920a, pp. 50–51). Online Appendix Figure A.3 shows the routes of the ferries. These descriptions illustrate Tsukishima as a busy machine manufacturing area during industrialization.
@@ 2.3 Risk-coping Institutions
Working-class households were exposed to several risks that led to losses of income at that time. Generally, social insurance, such as health, accident, and unemployment insurance, can mitigate idiosyncratic income shocks on workers. [^19] Gertler and Gruber (2002) shows the impact of illness on consumption-smoothing behavior in the early 1990s in Indonesia and argues for the importance of public compensation for illness. Kantor and Fishback (1996) show that the installation of workers’ compensation reduced the precautionary savings in the interwar US. Gruber (1997) provides evidence that unemployment insurance smooths individual consumption as was the case in the US from 1968–1987. However, comprehensive social insurance systems did not exist in the early 1920s in Japan. The previous health insurance was enacted in 1927, and unemployment insurance was not established throughout the prewar period. Additionally, the Factory Act enacted in 1916 did not specify compensation for the sickness of workers. Therefore, they had no public support for their illness, injuries, and unemployment. Instead, they might have used private financial institutions to manage idiosyncratic risks.
@@
Several financial institutions were available for working-class households. A representative savings institution was the postal savings (yūbin chokin). In Tokyo City, approximately $1.4$ million people had postal savings accounts, covering approximately $67$% of the citizens in 1920 (Online Appendix Table A.2). The average savings amount per capita in Kyobashi Ward was $44.5$ yen (Tokyo City Office 1921, pp. 892–893), which was less than one month’s earnings for the skilled-factory workers (Section 3.2). Savings bank (chochiku ginkō) was another famous savings institution among the working class in the 1910s (Ito and Saito 2019, pp. 79–81). The number of people with savings accounts was approximately $1.4$ million, accounting for approximately $58$% of total people in Tokyo (Online Appendix Table A.2). However, $84.8$% of all the depositors were commercial and miscellaneous workers, and the manufacturing sector shared only $11.5$% in Kyobashi Ward (Online Appendix Table A.3). This reflects the fact that savings banks were institutions designed for lower-income workers rather than factory workers (Tokyo Institute for Municipal Research 1925b, p. 81). This is consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices (Section 2.2). Although another savings institution was the mutual loan association (mujin), it was used among small business owners and merchants for running their businesses (Online Appendix A.2). [^20] Ordinary banks in prewar Japan provided large loans to businesses (Teranishi 2011). There was no bank in Tsukishima, as workers did not borrow money from banks to make ends meet. The official report states that “a commercial bank is not necessary in a laboring area such as Tsukishima, and even for large entrepreneurs such as factory owners, there is no particular need to establish a bank for their business transactions; transactions with banks in other areas are sufficient” (Department of Health, Ministry of the Interior 1923a, p. 48).
@@
There were also few lending institutions. Credit purchases (kakegai) with retailers were the most popular institution. Although systematic statistics of credit purchases are unavailable, the THBS households often used credit purchases in various retailers for rice, fish, vegetables, firewood, and charcoal. [^21] Cooperative societies were also available for worker households to buy daily commodities, although they were not lending institutions. There were $26$ cooperative societies in Tokyo prefecture in 1924, of which 23 were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p. 57–59). However, while the two largest societies had approximately $4,000$ members, the others were generally small societies with less than 100 members. This figure accounts for only a few percentage of the workers in the city, given that the number of male workers was more than $768$ thousand. In fact, the THBS households rarely used the cooperatives. I will describe this institution in detail in Section 5. The pawnshop (shichiya) was another lending institution. The accessibility was higher than for the other lending institutions. [^22] The number of pawnshops in Tokyo was $1,334$ and $1,261$ in 1918 and 1919, respectively. The total number of cases was $8,226,883$ and $7,573,406$ each year. These figures are more than three times greater than the entire population in Tokyo City. There were $13$ pawnshops in Tsukishima, more than twice the number of moneylenders (Section 2.2). However, the primary users of the pawnshops were the low-income working-class households rather than the factory workers (Tokyo City Social Affairs Bureau 1921, p. 9). According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the most representative users were the day laborers and workers classified as “miscellaneous,” accounting for $63.4$% of all users. By contrast, those for agriculture, commerce, and manufacturing sectors were only $1.7$, $19.5$, and $15.6$%, respectively (Tokyo Institute for Municipal Research 1926, p. 25). The inexpensive articles pawned suggest that the pawnshops were only used to dealing with very short-run loans for necessities. Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen, which accounts for $12$% of the monthly disposable income of the THBS households (Panel A of Table 3). Thus, the pawnshop was a small lending institution for low-income working-class households who needed credit. Indeed, lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the main articles pawned were inexpensive clothes (Shibuya et al. 1982). Additionally, the interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was substantially high under the lower interest rates. [^23] The average redemption rate was approximately $94$% in 1920 among $88$ pawnshops in Kyobashi ($401,222$ cases/$427,265$ cases) (Tokyo City Office 1922b, pp. 888–889).
Other lending institutions were money lenders (kinsen kashitsuke gyō) and credit unions (shinyō kumiai). Generally, moneylenders were used by business owners with very high interest rates (Shibuya 2000, pp. 184; 248). Similarly, the coverage of the credit unions was substantially low, possibly a few percentage points of the male workers in the manufacturing sector (Online Appendix A.2).
@@
Informal gifts might have been another type of risk-mitigation institution, although systematic statistics are not available. A household survey on the $185$ factory-worker households in Tokyo suggests that the share of average monthly gifts in total income was approximately $3$% in November 1922 (Social Affairs Division 1925, pp. 58–59). While this means that the gifts comprised a few percentage points of the monthly income, it is unclear whether these informal income transfers functioned as mutual aid.
@@
Finally, life insurance was not designed to compensate for temporary income losses. Generally, life insurance companies provided death insurance, meaning that temporary income reductions were not covered. Moreover, they did not target the working class as customers (Usami 1984, pp. 87–88). Although postal life insurance (kani seimei hoken) operated by the government started in 1916, it only provided life and endowment insurance. Thus, it never covered illness or unemployment (Usami 1984, pp. 102–107).  [^24] There were also several employment agencies that introduced jobs to low-income job-seekers in Tokyo. However, many of them were unscrupulous, and not commonly used at that time (Machida 2016).
@@
To summarize, while there were several lending institutions, few were available to the factory workers. Postal savings might have been the best available device for them. Credit purchase was the most plausible device regarding lending institutions. While pawnshops were accessible, they might not have provided sufficient money as the articles pawned were usually inexpensive. The factory workers in Tsukishima might have combined these devices to manage idiosyncratic income shocks.
@@ 3.1 Tsukishima Household Budget Survey
The Tsukishima Survey was the first urban social survey conducted in Japan by the Ministry of Home Affairs circa 1919. The purpose was to investigate the status of the lives of urban worker households amid the rapid industrialization throughout WWI. It was designed as a field survey consisting of multiple survey items, including the budget survey (hereinafter THBS). [^25] The THBS was the first household survey in Japan to use the budget book method and is considered the prototype for subsequent household surveys in Japan (Sekiya 1970, p. 43). Online Appendix B.1 summarizes detailed information about the Tsukishima Survey. The Ministry of Home Affairs published a few official reports on the Tsukishima Survey in 1921, which contained some aggregated information on the target households. However, household-level budget information is required to analyze consumption-smoothing strategies.
Fortunately, the household budget book (kinsendeiri hikaechō) gathered in the THBS is preserved in the archives of the Ohara Institute of Social Research (hereafter, OISR). [^26] The author was permitted to access those unreleased materials related to the Tsukishima Survey by the OISR: Archives of the Tsukishima Survey (THBS, unreleased). The official reports of the survey published in 1921 by the Ministry of Home Affairs are included in the archives of Gonda Yasunosuke at the OISR and are available to the public (The OISR, Archives of Gonda Yasunosuke (7-2; 7-3; 7-4)). These budget books were transferred to the OISR by Iwasaburo Takano, who was responsible for the Tsukishima Survey as a member of the 7th Section of the Health and Hygiene Research Committee of the Ministry of Home Affairs and became the first director of the OISR. [^27] Iwasaburo Takano was a professor at Tokyo Imperial University and a specialist in social statistics and social policy. He emphasized gathering statistics about workers through social surveys and is known as a researcher who significantly contributed to the formation of social survey methodology in Japan. His contribution to the Tsukishima Survey is summarized in detail in the Online Appendix B.1.
I corrected and digitized all forty THBS budget books for the factory worker households that were included in the official report. [^28] The archives of the OISR also preserve several other household budget books of factory worker households. However, these budget books contain poor record-keeping. Therefore, I do not use these incomplete books in this study. The characteristics of the remaining household budget books are summarized in detail in the Online Appendix B.5. During the study period, each household recorded their income and expenses daily in a budget book. Notably, a survey office was set up in Tsukishima, where a few surveyors lived and were closely connected with the survey households. Furthermore, weekly administrative meetings were held to discuss the progress of the survey (Miyoshi 1980, p. 38). This confirms that the information in these budget books is reliable and that the THBS budget dataset ensures a quality that enables quantitative analysis. [^29] Online Appendix B.1 provides finder details on how the Tsukishima Survey was conducted carefully as a social survey project of Takano.
@@ 3.2 Sample Characteristics
The estimation strategy used for testing consumption-smoothing and risk-coping strategies requires panel structures for aggregate income and expenditure (Section 4). I use the semi-monthly series for the baseline analysis, as the head’s income was paid twice a month. [^30] Section 3.3 provides details on the definitions of aggregate income and expenditure. Among the $40$ THBS households, there are five cross-sectional units in our semi-monthly data. I excluded these five cross-sectional units, leaving $35$ households. Next, I removed a household with incomplete income information. Finally, I excluded a household for which no information on payments to the credit purchases was recorded. Consequently, my THBS sample for the semi-monthly series included $33$ households with panel structures from January 12 to November 11, 1919.313131I also used the adjusted monthly panel dataset for sub-analysis in Section 4. Following the same trimming steps, my adjusted monthly panel dataset included $26$ households for the same period. Note that several units have only three semi-month cells, for example, less than two adjusted months. These units have a panel structure in the semi-monthly panel dataset, but not in the adjusted monthly panel. Online Appendix B.3 summarizes the finer details of the original data structure and trimming.
Before assessing the representatives of the THBS sample, I tested the potential selection bias due to attrition of the units. The panel units with shorter (longer) time-series observations might have different preferences on the consumption than those with longer (shorter) time-series observations. In other words, the households with shorter panels should have similar family size characteristics to those with more extended panels. To test this potential issue, I regressed an indicator variable for the shorter panel units on the family size variables. Online Appendix Table B.2 summarizes the results. All the estimated coefficients on the covariates were close to zero and statistically insignificant, and the Wald statistics support the null results. This is robust to using different thresholds such as first quantile, median, and third quantile. This supports the evidence that the household’s preference was less likely to be correlated with a lack of balance.
@@
# of workers $\geq 15$ workers $\geq 15$ workers $<15$
@@
Household size (share)
@@
Notes: Panel A: Column 1 shows the share of factory workers to the total factory workers in each sector (in percentage). The number of factory workers includes female and male factory workers in the 68 factories, with 15 workers and more. Column 2 indicates the share of factories to the total number of factories in each sector by factory size (in percentage). Column 3 summarizes the manufacturing sectors for the heads of the THBS. Columns 2 and 3 use the classification definition in the Statistics of the National Police Agency (column 1). Following this definition, several “woodworking” factories measured in the Tsukishima Factory Survey are included in the machinery sector in column 2, as it consists of shipbuilding and wooden pattern factories. The report also supports that the wooden patterns are used in the casting process and shall be included in the machinery sector (Department of Health, Ministry of the Interior 1923a, pp. 388–389).
Panel B: The first and second rows in columns 1 and 2 show the average household size and sex ratio measured in the 1920 Population Census and Statistics of the National Police Agency of 1919, respectively. The fourth to eighth rows in columns 1 and 3 show household size distributions in the census and the THBS sample, respectively. The figures in parentheses in Column (1) are the percentage share for family size bins between 2 and 8 people. The ninth row in column 3 shows the $p$-value from the Pearson $\chi^{2}$ test for the equality of the household size distributions between 2 and 8 people. The share in each household size bin measured in the census is used as the theoretical probability in calculating the test statistic (Online Appendix B.5).
Panel C: Column 1 lists the average monthly earnings of the THBS heads calculated using the adjusted monthly panel dataset. Of the 33 THBS households, information from 26 households for which adjusted monthly income can be traced back is used in this panel to follow the unit in the manufacturing census. Comparison based on the average semi-monthly income from all 33 households shows a similar result. Gaussian-based 95% confidence interval (CI) using bootstrap standard error is reported in the brackets. Column 2 shows the average monthly wage of the male factory workers in the skilled age range (30–40 years old) in machinery factories in Tokyo City. This figure is calculated using the average monthly wage and average monthly ancillary wage measured in the manufacturing censuses. All the wage statistics are deflated using the CPI for cities provided by Ohkawa et al. (1967, p. 255). Online Appendix B.6 summarizes the finer details of the process of this wage estimation.
Sources: Panel A: Column 1 uses the National Police Agency (1920b, p. 236). Column 2 uses the Department of Health, Ministry of the Interior (1923a, p. 385–386). Column 3 uses the Department of Health, Ministry of the Interior (1923a, p. 154). Panel B: Column 1 uses the Tokyo City Office (1922a, pp. 26–31; 262–283); Tokyo City Office (1922b, pp. 2–3; 42–46). Column 2 uses the National Police Agency (1920b, p. 53). The figures in column 3 are calculated by the author using the THBS dataset. Panel C: Column 1 uses the THBS dataset. Column 2 uses the Tokyo City Statistics Division (1926a, pp. 16; 244–245); Tokyo City Office (1921, pp. 726–757).
Next, I assess the representativeness of the THBS sample in Tsukishima using the manufacturing and population censuses.
@@
Panel A of Table 2 summarizes the share of male workers in the manufacturing industry by sector. Statistics of the National Police Agency (hereafter SNPA) covering factories with more than or equal to 15 workers show that approximately 90% of male workers were engaged in the machinery sector in 1919 (Column 1). Column 2 summarizes the statistics in the manufacturing census conducted in the Tsukishima Factory Survey. The proportion of machine factories is $85$% in larger factories and $92$% in smaller factories, indicating that most of the male workers worked in the machinery sector. [^32] I have confirmed that the paternal occupational statistics for the primary school students in Tsukishima provide materially similar results (Online Appendix A.5). In Column 3, I confirm that the THBS heads’ occupations show a similar proportion: $94$% worked in the machinery sector. In fact, the report states that “It is not too much to say that most of the industry in Tsukishima is machinery manufacturing. Therefore, to describe the labor situation in Tsukishima, it is sufficient to describe its machinery industry” (Department of Health, Ministry of the Interior 1923a, p. 389). Figure 3 illustrating a number of machinery factories supports this statement. As for the other sectors, the THBS includes a few heads who worked in food and miscellaneous sectors, and the proportions are similar to those listed in Column 2.
@@
Panel B of Table 2 summarizes the family structure. The average household sizes are identical in the Population Census (Column 1), SNPA (Column 2), and THBS (Column 3). The sex ratio is also similar across the surveys (1.21.21.21.21.31.31.31.3). The number of people in the THBS households ranges from 2 to 8, meaning the THBS sample does not include single or very large households. [^33] Those family households with 2–8 people cover most of the households in Tsukishima because the share of single and very large (9+ people) households was approximately $10$% (Column 1 of Panel B). In other words, the THBS mainly focuses on the consumption behavior in couples and households with a few children. For this household size range, the Pearson $\chi^{2}$ test does not reject the null of equality to the household size distribution measured in the census ($p$-value = $0.948$), meaning that the THBS sample reasonably approximates the distribution of family households in Tsukishima. Online Appendix B.5 illustrates the household size distributions in finer detail. This confirms the validity of the statement in the report that the THBS households can be “regarded as representative of the family form in Tsukishima” (Department of Health, Ministry of the Interior 1923a, p. 145).

### v14

@@ 2.2 Tsukishima: A Representative Manufacturing Area
Finally, I provide a brief overview of life in factory workers’ households. Factories and workers’ residences were scattered throughout Tsukishima. Factory workers lived in similar tenements and made use of a wide variety of commercial stores in the three lots: Tsukudajima and Shin-tsukudajima in the north, Lot 1 in the middle, and Lot 2 in the south (Figure 2). [^12] The northernmost Ishikawajima island contains only the Ishikawajima Shipyard, and the southernmost rectangular Lot 3 is uninhabited. Online Appendix Figures A.4 and C.5 provide detailed maps of the lots and retail shops, respectively. Online Appendix A.4 summarizes housing in Tsukishima. The main routes to the mainland were the wooden Aioi Bridge (Aioi bashi) and three ferries (watashi bune). The average daily number of ferry users in 1918 was $53,318$, more than twice the population of Tsukishima. [^13] Department of Health, Ministry of the Interior (1923a, p. 30). The number of people and households in Tsukishima at the end of December 1918 was $24,399$ and $5,562$, respectively (National Police Agency 1920a, pp. 50–51). Online Appendix Figure A.4 shows the ferry routes. These figures illustrate Tsukishima as a busy machine manufacturing area during industrialization.
@@ 2.3 Risk-coping Institutions
Generally, social insurance such as health, accident, and unemployment insurance can mitigate idiosyncratic income shocks (Gertler and Gruber 2002; Kantor and Fishback 1996; Gruber 1997). However, comprehensive social insurance systems did not exist in Japan in the early 1920s. Health insurance was not enacted until 1927, and unemployment insurance was absent throughout the prewar period. The Factory Act enacted in 1916 did not specify compensation for workers’ sickness. As a result, factory workers had no public support for illness, injuries, or unemployment.
@@
Instead, several financial institutions were available for factory workers. [^14] Finer details of these institutions and the statistics used in this subsection are summarized in Online Appendix A.2. A representative savings institution was postal savings (yūbin chokin). In Tokyo City, approximately $1.4$ million people held postal savings accounts, covering approximately 6767% of the citizens in 1920. The average savings per capita in Kyobashi Ward was $44.5$ yen, less than one month’s earnings for skilled factory workers (Section 3.2). Savings banks (chochiku ginkō) were another well-known savings institution. About $1.4$ million people held accounts, representing approximately 6767% of Tokyo’s population. However, 84.884.8% of all depositors were commercial and miscellaneous workers, while only 11.511.5% belonged to the manufacturing sector in Kyobashi Ward. This reflects that savings banks were designed for lower-income workers such as those in miscellaneous industries and day laborers, rather than factory workers (Tokyo Institute for Municipal Research 1925b, p. 81). This is also consistent with the fact that Tsukishima had no savings bank branches but did have a few post offices. Another savings institution, the mutual loan association (mujin), was mainly used by small business owners and merchants to run their businesses. [^15] Ordinary banks in prewar Japan provided large loans to businesses (Teranishi 2011). Tsukishima had no bank, since workers did not borrow money from banks to make ends meet (Department of Health, Ministry of the Interior 1923a, p. 48).
@@
Credit purchases (kakegai) with retailers were the most common informal credit institution. People often used credit purchases at various retailers for rice, fish, vegetables, firewood, and charcoal in Tokyo City. This institution is described in detail in Section 5. Pawnshops (shichiya) were widely distributed lending institutions in the city. [^16] The number of pawnshops in Tokyo was $1,334$ and $1,261$ in 1918 and 1919, with total transactions of $8,226,883$ and $7,573,406$ each year, respectively–more than three times the population of Tokyo City. However, evidence shows that the most frequent users were day laborers and workers classified as “miscellaneous,” accounting for 63.463.4% of all users. By contrast, the shares for agriculture, commerce, and manufacturing sectors were only $1.7$, $19.5$, and 15.615.6%, respectively (Tokyo Institute for Municipal Research 1926, p. 25). In practice, pawnshop lenders did not need to screen borrowers’ credit, and borrowers did not face the risk of heavy debt, as the main items pawned were inexpensive clothes (Shibuya et al. 1982). Interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was high under these lower rates. [^17] The average redemption rate was approximately 9494% in 1920 among $88$ pawnshops in Kyobashi (401,222401,222/$427,265$ cases) (Tokyo City Office 1922b, pp. 888–889). In short, the primary users of pawnshops were low-income workers who needed credit for daily necessities, rather than factory workers (Tokyo City Social Affairs Bureau 1921, p. 9). Other lending institutions included moneylenders (kinsen kashitsuke gyō) and credit unions (shinyō kumiai). Moneylenders were generally used by business owners and charged very high interest rates (Shibuya 2000, pp. 184; 248). Credit unions had very limited coverage, likely serving only a small share of male workers in the manufacturing sector.
@@
Informal gifts may have served as another type of risk-mitigation institution. While systematic statistics are unavailable, a household survey of $185$ factory-worker households in Tokyo found that the average monthly gifts accounted for only 33% of total income in November 1922 (Social Affairs Division 1925, pp. 58–59). Therefore, whether these small informal transfers functioned as effective mutual aid remains unclear.
@@
Life insurance was not designed to compensate for temporary income losses. Life insurance companies generally provided death insurance, which did not cover short-term income reductions, and they did not target working-class customers (Usami 1984, pp. 87–88). Although postal life insurance (kani seimei hoken) operated by the government began in 1916, it offered only life and endowment insurance and did not cover illness or unemployment (Usami 1984, pp. 102–107). [^18] Several employment agencies existed in Tokyo to introduce jobs to low-income job seekers. However, many were unscrupulous and not commonly used at the time (Machida 2016).
Finally, cooperative societies were available to help worker households purchase daily commodities. In 1924, Tokyo prefecture had $26$ cooperative societies, of which 23 were organized by citizens and workers (Central Federation of Industrial Associations 1925, p. 57–59). While the two largest societies had approximately $4,000$ members each, the others were generally small, with fewer than 100 members. Overall, this accounted for only a small share of workers in the city.
@@
To summarize, although several risk-coping institutions existed, few were accessible to factory workers. Postal savings were the main official financial institution available, while credit purchases were the most plausible informal mechanism. Factory workers in Tsukishima likely combined these devices to manage idiosyncratic income shocks.
@@ 3.1 Tsukishima Household Budget Survey
The Tsukishima Survey was the first urban social survey conducted in Japan by the Ministry of Home Affairs circa 1919. Its purpose was to examine the living conditions of urban worker households amid rapid industrialization during WWI. The survey was designed as a field study with multiple components, including a budget survey (hereinafter THBS). The Ministry of Home Affairs published several official reports in 1921, which contained aggregated information on the target households. [^19] The THBS was the first household survey to use the budget book method and is considered the prototype for subsequent household surveys in Japan (Sekiya 1970, p. 43). The official reports published in 1921 by the Ministry of Home Affairs are included in the archives of Gonda Yasunosuke and are available to the public (The OISR, Archives of Gonda Yasunosuke (7-2; 7-3; 7-4)). However, household-level budget data are necessary to analyze consumption-smoothing strategies.
Fortunately, the household budget books (kinsendeiri hikaechō) collected in the THBS are preserved in the archives of the Ohara Institute of Social Research (OISR). These budget books were transferred to the OISR by Professor Iwasaburo Takano of Tokyo Imperial University. He served on the research committee of the Ministry of Home Affairs responsible for the Tsukishima Survey and later became the first director of the OISR. [^20] Online Appendix B.1 provides further details on how this survey was conducted as part of Iwasaburo Takano’s ambitious social survey project.
I collected and digitized all forty THBS budget books for the factory-worker households. [^21] The author was permitted access to the unreleased materials related to the Tsukishima Survey by the OISR: Archives of the Tsukishima Survey (THBS, unreleased). The characteristics of the budget books are summarized in detail in Online Appendix B.2. During the study period, each household recorded its daily income and expenses in a budget book. Notably, a survey office was established in Tsukishima, where a few surveyors lived and maintained close contact with the surveyed households. Weekly administrative meetings were held to monitor the survey’s progress (Miyoshi 1980, p. 38).
@@ 3.2 Sample Characteristics
The estimation strategy requires panel structures for aggregate income and expenditure (Section 4). I use the semi-monthly series for the baseline analysis, as the household head’s income was paid twice a month. [^22] Section 3.3 provides details on the definitions of aggregate income and expenditure. Among the $40$ THBS households, five cross-sectional units appear in my semi-monthly data. I exclude these, leaving $35$ households. I then remove a household with incomplete income information and another for which no information on payments for credit purchases are recorded. Consequently, the THBS sample for the semi-monthly series includes $33$ households with panel structures from January 12 to November 11, 1919.232323I also use the adjusted monthly panel dataset for the sub-analysis in Section 4. Following the same trimming steps, the adjusted monthly panel dataset includes $26$ households for the same period. Note that several units have only three semi-month cells, for example, less than two adjusted months. These units have a panel structure in the semi-monthly dataset but not in the adjusted monthly panel. Online Appendix B.3 summarizes the original data structure and trimming in detail.
Before assessing the representativeness of the THBS sample, I evaluate potential selection bias due to unit attrition. Panels with shorter (or longer) time-series observations might have different consumption preferences than those with longer (or shorter) observations. In other words, households with shorter panels should have family size characteristics similar to those with longer panels. To test this, I regress an indicator variable for the shorter panel units on family size variables. Online Appendix Table B.2 summarizes the results. All estimated coefficients on the covariates are close to zero and statistically insignificant, and the Wald statistics support the null hypothesis. This result is robust to using different thresholds such as the first quantile, median, and third quantile. These findings suggest that household preferences are unlikely to be correlated with panel length.
@@
# of workers $\geq 15$ workers $\geq 15$ workers <15<15
@@
Household size (% share)
@@
Notes: Panel A: Column 1 shows the share of factory workers in each sector relative to the total number of factory workers (percentage). This includes both female and male workers in the 68 factories with 15 or more employees. Column 2 indicates the share of factories in each sector by factory size (percentage). Column 3 summarizes the manufacturing sectors of the THBS household heads. Columns 2 and 3 follow the classification used in the Statistics of the National Police Agency (Column 1). According to this classification, several “woodworking” factories recorded in the Tsukishima Factory Survey are included in the machinery sector in Column 2, as it covers shipbuilding and wooden pattern factories. The report also confirms that wooden patterns are used in the casting process and should therefore be included in the machinery sector (Department of Health, Ministry of the Interior 1923a, pp. 388–389). Panel B: The first and second rows in Columns 1 and 2 show the average household size and sex ratio from the 1920 Population Census and the 1919 Statistics of the National Police Agency, respectively. Rows four to eight in Columns 1 and 3 present household size distributions in the census and the THBS sample, respectively. Figures in parentheses in Column 1 indicate the percentage share for family size bins from 2 to 8 people. Row nine in Column 3 reports the $p$-value from the Pearson $\chi^{2}$ test for equality of household size distributions between 2 and 8 people. The census shares in each household size bin are used as the theoretical probabilities in calculating the test statistic (Online Appendix B.5). Panel C: Column 1 lists the average monthly earnings of THBS household heads, calculated from the adjusted monthly panel dataset. Of the 33 THBS households, data from 26 households–where adjusted monthly income can be traced–are used in this panel to match units in the manufacturing census. Comparison using average semi-monthly income from all 33 households produces similar results. Gaussian-based 95% confidence intervals (CI) using bootstrap standard errors are reported in brackets. Column 2 shows the average monthly wage of male factory workers aged 30–40 in machinery factories in Tokyo City. This figure is calculated using average monthly wages and ancillary wages from the manufacturing censuses. All wage statistics are deflated using the CPI for cities from Ohkawa et al. (1967, p. 255). Online Appendix B.6 provides further details on the wage estimation process. Sources: Panel A: Column 1: National Police Agency (1920b, p. 236); Column 2: Department of Health, Ministry of the Interior (1923a, p. 385–386); Column 3: Department of Health, Ministry of the Interior (1923a, p. 154). Panel B: Column 1: Tokyo City Office (1922a, pp. 26–31; 262–283); Tokyo City Office (1922b, pp. 2–3; 42–46); Column 2: National Police Agency (1920b, p. 53); Column 3: Calculated by the author from the THBS dataset. Panel C: Column 1: THBS dataset; Column 2: Tokyo City Statistics Division (1926a, pp. 16; 244–245); Tokyo City Office (1921, pp. 726–757).
@@
Panel A of Table 2 summarizes the share of male workers in the manufacturing industry by sector. Statistics from the National Police Agency (SNPA) for factories with 15 or more workers show that approximately 9090% of male workers were employed in the machinery sector in 1919 (Column 1). Column 2 presents statistics from the Tsukishima Factory Survey. The proportion of machinery factories is 8585% among larger factories and 9292% among smaller factories, indicating that most male workers were in the machinery sector. [^24] Paternal occupational statistics for primary school students in Tsukishima yield materially similar results (Online Appendix A.5). Column 3 confirms that THBS household heads show a similar pattern, with 9494% working in the machinery sector. The report states: “It is not too much to say that most of the industry in Tsukishima is machinery manufacturing. Therefore, to describe the labor situation in Tsukishima, it is sufficient to describe its machinery industry” (Department of Health, Ministry of the Interior 1923a, p. 389). The rest of the THBS heads work in food or miscellaneous industries, with proportions similar to those in Column 2.
@@
Panel B of Table 2 summarizes family structure. Average household sizes are identical across the Population Census (Column 1), SNPA (Column 2), and THBS (Column 3). Sex ratios are also similar across the surveys (1.21.21.31.3). THBS households range from 2 to 8 members, excluding single-person and very large households. In other words, the THBS mainly captures consumption behavior in couples and households with a few children. For this range, the Pearson $\chi^{2}$ test does not reject the null hypothesis of equality with the census household size distribution ($p$-value = $0.948$), indicating that the THBS sample reasonably approximates the distribution of family households in Tsukishima. Online Appendix B.5 provides a more detailed view of household size distributions. This supports the report’s statement that THBS households can be “regarded as representative of the family form in Tsukishima” (Department of Health, Ministry of the Interior 1923a, p. 145).